What value of m satisfies the equation shown below?
5(m - 5) = 3(m + 1)

Answer by checkley79(3341)

5(m - 5) = 3(m + 1)
5m-25=3m+3
5m-3m=3+25
2m=28
m=28/2
m=14 ans.